OpenStudy (anonymous):

Evaluate \[\sum_{k=0}^{n}(-4)^{k}\](n+k choose 2k) in closed form.

OpenStudy (anonymous):

it should be mulitiplied...

OpenStudy (anonymous):

did that help

OpenStudy (anonymous):

I'm going to check it and then let you know... thanks

OpenStudy (jamesj):

You want \[ \sum_{k=0}^n (-4)^k { n+k \choose 2k } \] yes?

OpenStudy (anonymous):

yes

OpenStudy (jamesj):

Right. Let's test a few values and see if we can detect a pattern n=0: sum = 1 n=1: \[ \Sigma = (-4)^0 { 0 \choose 0} + (-4)^1 { 2 \choose 2} = 1 - 4 = -3 \] n=2: \[ \Sigma = (-4)^0 { 2 \choose 0} + (-4)^1 { 3 \choose 2} + (-4)^2 { 4 \choose 4} \] \[ = 1 -4(3) + 16(1) = 5 \]

OpenStudy (jamesj):

If you calculate this further, you'll find n = 3, sum = -7 n = 4, sum = 9 n = 5, sum = -11

OpenStudy (jamesj):

Hence it looks like the sum is equal to \[ (-1)^n(2n+1) \] To prove that, now, use induction and/or other identities you have for binomial coefficients.

OpenStudy (anonymous):

Thanks, we were just confused with what the sum would be equal to...I'll get to work on that